The table below shows the time in hours (𝑥) spent by six (6) students in playing


Clash of Clansand the scores of these students got on a math test. Solve for


Pearson Product Correlation Coeffcient. Construct a Scatter Plot.


𝑥 1 2 3 4 5 6


𝑦 30 25 25 10 15 5

1
Expert's answer
2022-06-09T14:59:06-0400

In order to compute the regression coefficients, the following table needs to be used:


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
 & X & Y & XY & X^2 & Y^2 \\ \hline
 & 1 & 30 & 30 & 1 & 900 \\
 \hdashline
 & 2 & 25 & 50 & 4 & 625 \\
 \hdashline
 & 3 & 25 & 75 & 9 & 625 \\
 \hdashline
 & 4 & 10 & 40 & 16 & 100 \\
 \hdashline
 & 5 & 15 & 75 & 25 & 225 \\
 \hdashline
 & 6 & 5 & 30 & 36 & 25 \\
 \hdashline
Sum= & 21 & 110 & 300 & 91 & 2500 \\
 \hdashline
\end{array}""\bar{X}=\dfrac{1}{n}\sum _{i}X_i=\dfrac{21}{6}=\dfrac{7}{2}"




"\bar{Y}=\dfrac{1}{n}\sum _{i}Y_i=\dfrac{110}{6}=\dfrac{55}{3}"




"SS_{XX}=\sum_iX_i^2-\dfrac{1}{n}(\sum _{i}X_i)^2""=91-\dfrac{21^2}{6}=\dfrac{35}{2}"




"SS_{YY}=\sum_iY_i^2-\dfrac{1}{n}(\sum _{i}Y_i)^2""=2500-\dfrac{110^2}{6}=\dfrac{1450}{3}"




"SS_{XY}=\sum_iX_iY_i-\dfrac{1}{n}(\sum _{i}X_i)(\sum _{i}Y_i)""=300-\dfrac{21(110)}{6}=-85"




"r=\dfrac{SS_{XY}}{\sqrt{SS_{XX}SS_{YY}}}=\dfrac{-85}{\sqrt{\dfrac{35}{2}(\dfrac{1450}{3})}}""=-0.9242"


Strong negative correlation.